# On the Jacobs-de Leeuw-Glicksberg decomposition

### v1

@@ 2.4 Representations of semigroups
Let $E$ be a Banach space and $S$ a semigroup. A representation $\pi$ of $S$ on $E$ is a weakly continuous [^5] We warn the reader that many other texts require that $\pi$ be strongly continuous. Nonetheless, for our purposes weak continuity is enough. map $\pi:S\rightarrow\mathcal{L}(E)$. If $S$ is a monoid with identity $e$, then we will assume that $\pi_{e}$ is the identity map. We note that if $\mathcal{H}$ is a Hilbert space and $U$ is a unitary representation of $S$ on $\mathcal{H}$, then $U$ is strongly continuous. The representation $\pi$ is uniformly bounded if $\sup_{s\in S}\|\pi_{s}\|<\infty$, and it is relatively weakly compact if $\{\pi_{s}\ |\ s\in S\}$ is relatively compact in the weak topology of $\mathcal{L}(E)$. A fact that we will frequently use is that if $\pi$ is a relatively weakly compact representation of $S$ on $E$, then for any $\xi\in E$ and any $f^{\prime}\in E^{\prime}$, the function $\phi(s)=f^{\prime}(\pi_{s}\xi)$ belongs to $W(S)$ (cf. [17, Theorem 2.7]). While every relatively weakly compact set of operators in $\mathcal{L}(E)$ is uniformly bounded, the converse need not be true.
@@
A matrix coefficient $\phi$ of the representation $\pi$ is a function of the form $\phi(s)=f^{\prime}(\pi_{s}\xi)$ for some $\xi\in E$ and some $f^{\prime}\in E^{\prime}$. If $\pi_{1}$ and $\pi_{2}$ are representations of $S$ on $E_{1}$ and $E_{2}$ respectively, then $\pi_{1}$ is weakly contained in $\pi_{2}$ if every matrix coefficient of $\pi_{1}$ can be approximated arbitrary well by linear combinations of matrix coefficients of $\pi_{2}$. To be more precise, $\pi_{1}$ is weakly contained in $\pi_{2}$ if for any $\xi\in E_{1}$, any $f^{\prime}\in E_{1}^{\prime}$, and any $\epsilon>0$, there exists $\eta_{1},\cdots,\eta_{k}\in E_{2}$ and $h^{\prime}_{1},\cdots,h^{\prime}_{k}\in E^{\prime}_{2}$ for which
@@
The representations $\pi_{1}$ and $\pi_{2}$ are weakly equivalent if $\pi_{1}$ is weakly contained in $\pi_{2}$ and $\pi_{2}$ is weakly contained in $\pi_{1}$. The representations $\pi_{1}$ and $\pi_{2}$ are equivalent if there exists a bounded linear invertible operator $T:E_{1}\rightarrow E_{2}$ for which $\pi_{2,s}T=T\pi_{1,s}$ for all $s\in S$. It is known that if $T:E_{1}\rightarrow E_{2}$ is a bounded linear invertible map, then $E_{1}$ is a Hilbert space if and only if $E_{2}$ is a Hilbert space, so a representation of $S$ on a Banach space $E$ that is not a Hilbert space cannot be equivalent to any representation of $S$ on a Hilbert space $\mathcal{H}$. However, as we shall see in Theorem 3.4, there are many representations of $S$ on a Banach space that are weakly equivalent to a representation on a Hilbert space.
@@ 2.5 Ultrafilters and the algebra of the Stone-Čech compactification
We embed $S$ into $\beta S$ as a dense subset by identifying each $s\in S$ with the ultrafilter $e(s):=\{A\subseteq S|s\in A\}$. When $(S,\cdot)$ is a discrete semigroup, there is a unique extension of the operation $\cdot$ from $e(S)$ to all of $\beta S$ that makes $\beta S$ a compact right topological semigroup, i.e., each of the maps $R_{p}$ for $p\in\beta S$ are continuous. One can just as well extend the semigroup operation from $S$ to $\beta S$ so as to make each of the maps $L_{p}$ continuous instead of $R_{p}$, but in general the operation cannot be extended in such a way as to make $\beta S$ a semitopological semigroup. If $\mathscr{S}$ is a compact semitopological semigroup and $i:S\rightarrow\mathscr{S}$ is a continuous homomorphism whose image is dense in $\mathscr{S}$, then there exists a unique continuous surjective homomorphism $h:\beta S\rightarrow\mathscr{S}$ satisfying $h\circ e=i$, and it is given by $h(p)=p-\lim_{s}e(s)$ (cf. [29, Theorem 4.8]). It is known that $\beta S$ possesses a smallest two-sided ideal that we denote by $K(\beta S)$, and that $K(\beta S)$ possesses idempotents.
If $S$ is a left-amenable [^7] We mention that we work with a left-amenable semigroup instead of a right-amenable semigroup because we made $\beta S$ a right topological semigroup. semigroup satisfying the strong Følner condition, then we let $\triangle^{*}(S)=\{p\in\beta S|\forall\ A\in p,d^{*}(A)>0\}$. In [26, Theorem 2.8] it is shown that $\triangle^{*}(S)$ is a closed two-sided ideal in $\beta S$, hence it contains $\overline{K(\beta S)}$. An idempotent element of $\triangle^{*}(S)$ is called an essential idempotent. Essential idempotents play an important role in ergodic Ramsey theory (cf. [10, 35, 6, 11, 15]).
For further discussion about the algebra of $\beta S$ when $S$ is a discrete semigroup, the reader is referred to [29]. When mention that the algebra of Stone-Čech compactification of a semitopological semigroup $S$ has been investigated in [2], but the situation there is much more complicated, which is why we choose to restrict our attention to discrete semigroups when working with the Stone-Čech compactification.
@@ 3 The reversible part of the JdLG-decomposition
As in section 2.2 we denote by $\mathscr{T}$ a relatively weakly compact operator semigroup acting on a Banach space $E$ and we denote by $\mathscr{S}$ the weak closure of $\mathscr{T}$. We have already seen that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a (compact topological) group. In this case there exists a unique minimal idempotent $Q\in\mathscr{S}$. The reversible part of $E$ is defined by setting $E_{r}:=\operatorname{ran}(Q)$. Letting $\mathscr{S}_{r}$ denote the restriction of $\mathscr{S}$ to $E_{r}$, it is known that $\mathscr{S}_{r}$ is a compact topological group that is (algebraically and topologically) isomorphic to $K(\mathscr{S})$ (cf. [21, Theorem 16.24]). We call a tuple $(e_{j})_{j=1}^{n}\in E^{n}$ a finite unitary system for $\mathscr{T}$ if the vectors $e_{1},...,e_{n}$ are linearly independent, the set $F:=\operatorname{lin}\{e_{1},...,e_{n}\}$ is $\mathscr{T}$-invariant, and the corresponding matrix representation $\chi$ of $\mathscr{T}$ defined by
@@
is unitary, i.e. for every $T\in\mathscr{T}$, $\chi(T)$ is an unitary $n\times n$ matrix. Moreover an unitary system $(e_{j})_{j=1}^{n}$ is called irreducible if $F=\operatorname{lin}\{e_{1},...,e_{n}\}$ does not contain any nontrivial $\mathscr{T}$-invariant subspaces. By [21], Theorem 16.31 the reversible part of a JdLG-admissable semigroup $\mathscr{T}$ is given by
@@
Under an additional assumption we can characterize the reversible part as follows.
@@
Since $\pi$ is JdLG-admissible, it is weakly almost periodic, so there exists $C>0$ such that $\sup_{g\in G}\|\pi_{g}\|\leq C$. Hence, for any $x\in E$ we obtain
@@
Let $S$ be a semigroup, $E$ a (Banach) Hilbert space, $\pi$ a JdLG-admissible representation of $S$ on $E$, and $\pi_{r}$ the restriction of $\pi$ to $E_{r}$. The representation $\pi_{r}$ is (weakly) equivalent to a unitary representaiton $U$ of $S$ on a Hilbert space $\mathcal{H}$.
@@
Let us first assume that $E$ is a general Banach space. Let $(F_{i})_{i\in I}$ denote the collection of finite dimensional subspaces of $E_{r}$ that have an irreducible finite unitary systems $\{e_{i,j}\}_{j=1}^{n_{i}}$ as a basis. We see that for $i\neq j\in I$, $F_{i}\cap F_{j}$ is a $\pi(S)$-invariant subspace of $F_{i}$, hence $F_{i}\cap F_{j}=\{0\}$. For $i\in I$ and $1\leq j\leq n_{i}$ let $f_{i,j}^{\prime}\in E_{r}^{\prime}$ be the functional that is 0 on the complement of $F_{i}$ and on $F_{i}$ satisfies
$f_{i,j}^{\prime}\left(\sum_{k=1}^{n_{i}}c_{i,k}e_{i,k}\right)=c_{i,j}.$ (4)
Since $E_{r}$ is the smallest closed subspace containing all of the $F_{i}$, we see that the closed linear hull of $\{\{e_{i,j}\}_{j=1}^{n_{i}}\}_{i\in I}$ is $E_{r}$, and that the closed linear hull of $\{\{f_{i,j}^{\prime}\}_{j=1}^{n_{i}}\}_{i\in I}$ is $E_{r}^{\prime}$. Let $\mathcal{H}_{i}=\mathbb{C}^{n_{i}}$ for $i\in I$. For $s\in S$ and $i\in I$, let $\pi_{i,s}$ denote the $n_{i}\times n_{i}$ unitary matrix that represents the action of $\pi_{s}$ on $F_{i}$ with respect to the basis $\{e_{i,j}\}_{j=1}^{n_{i}}$. Let $U_{i,s}$ denote the same matrix as $\pi_{i,s}$, but acting on $\mathbb{C}^{n_{i}}$ with the standard basis instead of $F_{i}$. It follows that $U=\oplus_{i\in I}U_{i}$ is a unitary representation of $S$ acting on the Hilbert space $\mathcal{H}:=\oplus_{i\in I}\mathbb{C}^{n_{i}}$. We let $\phi:E_{r}\rightarrow\mathcal{H}$ be the map that identifies $\{\{e_{i,j}\}_{j=1}^{n_{i}}\}_{i\in I}$ with the canonical orthonormal basis of $\mathcal{H}$, and we observe that $\phi$ is not necessarily a bounded linear invertible map. For $\xi\in\mathcal{H}$, let $\xi^{\prime}\in\mathcal{H}^{\prime}$ be the functional given by $\xi^{\prime}(\eta)=\langle\eta,\xi\rangle$. To see that the representation $U$ is weakly equivalent to $\pi_{r}$, it suffices to observe that for any $i_{1},i_{2}\in I$ and $(j_{1},j_{2})\in[1,n_{i_{1}}]\times[1,n_{i_{2}}]$, we have $\phi(e_{i_{1},j_{1}})^{\prime}(\phi(e_{i_{2},j_{2}}))=f_{i_{1},j_{1}}^{\prime}(e_{i_{2},j_{2}})$.
Now let us consider the case in which $E$ is a Hilbert space. We see that $E=\oplus_{i\in I}F_{i}$. We recall that $\mathscr{S}_{r}:=\overline{\pi_{r}(S)}$ is a compact (hence amenable) topological group. Day [16] showed that any uniformly bounded representation of an amenable group $G$ on a Hilbert space is unitarizable, i.e., is equivalent to a unitary representation of $G$. The identity map gives a uniformly bounded representation of $\mathscr{S}_{r}$ on the Hilbert space $E$, so it is unitarizable. Since $E$ decomposes into a direct sum of finite dimensional representations of $\mathscr{S}_{r}$, any equivalent representation does the same, which yields the desired result. ∎

### v2

@@ 2.4 Representations of semigroups
Let $E$ be a Banach space, $\mathcal{L}(E)$ the space of bounded linear operators on $E$, and $S$ a semigroup. A representation $\pi$ of $S$ on $E$ is a weakly continuous [^6] We warn the reader that many other texts require that $\pi$ be strongly continuous. Nonetheless, for our purposes weak continuity is enough. map $\pi:S\rightarrow\mathcal{L}(E)$. If $S$ is a monoid with identity $e$, then we will assume that $\pi_{e}$ is the identity map. We note that if $\mathcal{H}$ is a Hilbert space and $U$ is a unitary representation of $S$ on $\mathcal{H}$, then $U$ is strongly continuous. The representation $\pi$ is uniformly bounded if $\sup_{s\in S}\|\pi_{s}\|<\infty$, and it is relatively weakly compact if $\{\pi_{s}\ :\ s\in S\}$ is relatively compact in the weak topology of $\mathcal{L}(E)$. A fact that we will frequently use is that if $\pi$ is a relatively weakly compact representation of $S$ on $E$, then for any $\xi\in E$ and any $f^{\prime}\in E^{\prime}$, the function $\phi(s)=f^{\prime}(\pi_{s}\xi)$ belongs to $W(S)$ (cf. [16, Theorem 2.7]). Since $W(S)$ is a $C^{*}$-algebra, it follows that $|\phi|^{p}\in W(S)$ as well for every $p>0$. While every relatively weakly compact set of operators in $\mathcal{L}(E)$ is uniformly bounded, the converse need not be true.
@@
A matrix coefficient $\phi$ of the representation $\pi$ is a function of the form $\phi(s)=f^{\prime}(\pi_{s}\xi)$ for some $\xi\in E$ and some $f^{\prime}\in E^{\prime}$. If $\pi_{1}$ and $\pi_{2}$ are representations of $S$ on $E_{1}$ and $E_{2}$ respectively, then $\pi_{1}$ is weakly contained in $\pi_{2}$ if every matrix coefficient of $\pi_{1}$ can be approximated arbitrary well by linear combinations of matrix coefficients of $\pi_{2}$. To be more precise, $\pi_{1}$ is weakly contained in $\pi_{2}$ if for any $\xi\in E_{1}$, any $f^{\prime}\in E_{1}^{\prime}$, and any $\epsilon>0$, there exist $\eta_{1},\cdots,\eta_{k}\in E_{2}$ and $h^{\prime}_{1},\cdots,h^{\prime}_{k}\in E^{\prime}_{2}$ for which
@@
The representations $\pi_{1}$ and $\pi_{2}$ are weakly equivalent if $\pi_{1}$ is weakly contained in $\pi_{2}$ and $\pi_{2}$ is weakly contained in $\pi_{1}$. The representations $\pi_{1}$ and $\pi_{2}$ are equivalent if there exists a bounded linear invertible operator $T:E_{1}\rightarrow E_{2}$ for which $\pi_{2,s}T=T\pi_{1,s}$ for all $s\in S$.
@@ 2.5 Ultrafilters and the algebra of the Stone-Čech compactification
We embed $S$ into $\beta S$ as a dense subset by identifying each $s\in S$ with the ultrafilter $e(s):=\{A\subseteq S|s\in A\}$. When $(S,\cdot)$ is a discrete semigroup, there is a unique extension of the operation $\cdot$ from $e(S)$ to all of $\beta S$ that makes $\beta S$ a compact right topological semigroup, i.e., each of the maps $R_{p}$ for $p\in\beta S$ are continuous. One can just as well extend the semigroup operation from $S$ to $\beta S$ so as to make each of the maps $L_{p}$ continuous instead of $R_{p}$, but in general the operation cannot be extended in such a way as to make $\beta S$ a semitopological semigroup. If $\mathscr{S}$ is a compact semitopological semigroup and $i:S\rightarrow\mathscr{S}$ is a continuous homomorphism whose image is dense in $\mathscr{S}$, then there exists a unique continuous surjective homomorphism $h:\beta S\rightarrow\mathscr{S}$ satisfying $h\circ e=i$, and it is given by $h(p)=p-\lim_{s}e(s)$ (cf. [27, Theorem 4.8]). It is known that $\beta S$ possesses the smallest two-sided ideal that we denote by $K(\beta S)$, and that $K(\beta S)$ possesses idempotents.
If $S$ is a left-amenable [^8] We mention that we work with a left-amenable semigroup instead of a right-amenable semigroup because we made $\beta S$ a right topological semigroup. semigroup satisfying the strong Følner condition, then we let $\triangle^{*}(S)=\{p\in\beta S|\forall\ A\in p,d^{*}(A)>0\}$. In [25, Theorem 2.8] it is shown that $\triangle^{*}(S)$ is a closed two-sided ideal in $\beta S$, hence it contains $\overline{K(\beta S)}$. An idempotent element of $\triangle^{*}(S)$ is called an essential idempotent. Essential idempotents play an important role in ergodic Ramsey theory (cf. [9, 34, 6, 10, 14]).
For further discussion about the algebra of $\beta S$ when $S$ is a discrete semigroup, the reader is referred to [27]. We mention that the algebra of Stone-Čech compactification of a semitopological semigroup $S$ has been investigated in [2], but the situation there is much more complicated, which is why we choose to restrict our attention to discrete semigroups when working with the Stone-Čech compactification.
@@ 3 The reversible part of the JdLG-decomposition
As in section 2.2, we denote by $\mathscr{T}$ a relatively weakly compact operator semigroup acting on a Banach space $E$, and we denote by $\mathscr{S}$ the weak closure of $\mathscr{T}$. We have already seen that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a (compact topological) group. In this case there exists a unique minimal idempotent $Q\in\mathscr{S}$. The reversible part of $E$ is defined by setting $E_{r}:=\operatorname{ran}(Q)$, and as discussed above, it is also described by (1). Letting $\mathscr{S}_{r}$ denote the restriction of $\mathscr{S}$ to $E_{r}$, it is known that $\mathscr{S}_{r}$ is a compact topological group that is (algebraically and topologically) isomorphic to $K(\mathscr{S})$ (cf. [20, Theorem 16.24]). We call a tuple $(e_{j})_{j=1}^{n}\in E^{n}$ a finite unitary system for $\mathscr{T}$ if the vectors $e_{1},...,e_{n}$ are linearly independent, the set $F:=\operatorname{lin}\{e_{1},...,e_{n}\}$ is $\mathscr{T}$-invariant, and the corresponding matrix representation $\chi$ of $\mathscr{T}$ defined by
@@
is unitary, i.e. for every $T\in\mathscr{T}$, $\chi(T)$ is an unitary $n\times n$ matrix. If $(e_{j})_{j=1}^{n}$ is a finite unitary system for $\mathscr{T}$ with the corresponding matrix representation $\chi$, then for each $1\leq i,j\leq n$ the map $T\mapsto\chi_{ij}(T)$ is continuous. A unitary system $(e_{j})_{j=1}^{n}$ is called irreducible if $F=\operatorname{lin}\{e_{1},...,e_{n}\}$ does not contain any nontrivial $\mathscr{T}$-invariant subspaces. By [20], Theorem 16.31 the reversible part of a JdLG-admissable semigroup $\mathscr{T}$ is given by
@@
Under an additional assumption one can characterize the reversible part as follows.
@@
Since $\pi$ is JdLG-admissible, it is weakly relatively compact, thus uniformly bounded, so there exists $C>0$ such that $\sup_{g\in G}\|\pi_{g}\|\leq C$. Hence, for any $x\in E$ we obtain
@@
The following proposition presents another particular case of JdLG-admissible semigroups.
Proposition 3.4.
Let $E$ be a Banach space and let $\mathscr{S}$ be a weakly compact semigroup of bounded linear operators on $E$ such that every $\mathscr{S}$-orbit contains a unique fixed point. Then $\mathscr{S}$ is JdLG-admissible. Moreover, $E_{r}$ is the set of fixed vectors of $\mathscr{S}$.
Proof.
Let $E_{1}$ denote the set of fixed points of $\mathscr{S}$, which is a closed linear subspace of $E$. For any $\xi\in E$ let $\xi_{1}=u_{1}\xi$ be the fixed point in $\mathscr{S}\xi$. Then $u_{1}(\xi-\xi_{1})=0$, so $\xi_{0}:=\xi-\xi_{1}\in E_{\text{aws}}$, which shows that $E=E_{1}+E_{\text{aws}}$. The set $E_{\text{aws}}$ is also a linear subspace: if $\xi_{1},\xi_{2}\in E_{\text{aws}}$ and $u_{1}\xi_{1}=0$, then $u_{1}(\xi_{1}+\xi_{2})=u_{1}\xi_{2}$; by assumption, $\mathscr{S}u_{1}\xi_{2}$ contains a fixed point $vu_{1}\xi_{2}$, but it will also be a fixed point in $\mathscr{S}\xi_{2}$ which is 0, so that $vu_{1}(\xi_{1}+\xi_{2})=0$ and $\xi_{1}+\xi_{2}\in E_{\text{aws}}$. (Stability under multiplication by scalars is obvious.) The intersection of these subspaces is clearly 0, so we have a direct sum decomposition.
The space $E_{1}$ is in fact $E_{r}$ as defined in Theorem 1.1: if $\xi\in E_{r}$ and $\xi=\xi_{1}+\xi_{0}$ is its decomposition as above, then choose $u\in\mathscr{S}$ such that $u\xi_{0}=0$. For any $v\in\mathscr{S}$ we have then $vu\xi=\xi_{1}$, so choosing $v$ such that $vu\xi=\xi$ we obtain $\xi=\xi_{1}$ and thus $E_{r}\subset E_{1}$. The converse inclusion is obvious: (1) holds with any $u,v\in\mathscr{S}$. ∎
Example 3.5.
If the fixed point is not unique, the statement may not hold. Consider the semigroup of two elements $S=\{a,b\}$ with multiplication $a^{2}=ba=a$, $b^{2}=ab=b$. It acts on $E=C(S)$ by right translations, so that $R_{a}f\equiv f(a)$ and $R_{b}f\equiv f(b)$ for every $f$. A function is fixed under this action if and only if it is constant, and is in the set $E_{\text{aws}}$ if it vanishes at $a$ or at $b$. But this set is not a linear subspace, and the decomposition in $E_{r}+E_{\text{aws}}$, though possible, is not unique.
Proposition 3.4 applies to the following special semigroups not covered, in general, by Theorem 1.1: $\mathscr{S}$ is the weak closure of $\pi(P)$ where $P$ is the semigroup of normal states of a quantum semigroup $M$, and $\pi$ its so called $\ell_{2}$-bounded representation on a reflexive Banach space (see the proof of Proposition 3.1 in [33]). One example of such a semigroup $P$ is the set of probability measures on a locally compact group $G$, with convolution as the semigroup operation. This particular semigroup is known to have a two-sided invariant mean on $W(S)$ and thus falls into assumptions of Theorem 1.1, but in general this is not the case.
@@
Let $S$ be a semigroup, $E$ a (Banach) Hilbert space, $\pi$ a JdLG-admissible representation of $S$ on $E$, and $\pi_{r}$ the restriction of $\pi$ to $E_{r}$. The representation $\pi_{r}$ is (weakly) equivalent to a unitary representation $U$ of $S$ on a Hilbert space $\mathcal{H}$.
@@
Let us first consider the case in which $E$ is a Banach space. Let $X_{n}$ denote the set of matrix representations $\chi$ of $\pi(S)$ corresponding to some unitary system $(e_{j})_{j=1}^{n}$ in $E_{r}$, and let $X=\bigcup_{n=1}^{\infty}X_{n}$. For each $n\in\mathbb{N}$ and $\chi\in X_{n}$, let $\mathcal{H}_{\chi}=\mathbb{C}^{n}$, and observe that $\tilde{\chi}:=\chi\circ\pi$ can be viewed as a continuous representation of $S$ on $\mathcal{H}_{\chi}$. Let $\mathcal{H}=\displaystyle\oplus_{\chi\in X}\mathcal{H}_{\chi}$ be the Hilbert space direct sum of these spaces, and observe that $U=\oplus_{\chi\in X}\tilde{\chi}$ is a unitary representation of $S$ on $\mathcal{H}$ that decomposes into a direct sum of finite dimensional representations.
We now show that $U$ is weakly contained in $\pi_{r}$. Every coefficient $\phi(s)=f^{\prime}(U_{s}\xi)$ of $U$ is approximated (uniformly on $S$) by coefficients with $\xi$ supported only in a finite number of $\mathcal{H}_{\chi}$, and by linearity, it is sufficient to consider only $\phi$ such that $\xi\in\mathcal{H}_{\chi}$ with a single $\chi\in X$. Fix $\chi\in X$, and let $(e_{j})_{j=1}^{n}$ be a corresponding unitary system for $\pi_{r}(S)$. For $F=\text{lin}\{e_{1},\cdots,e_{n}\}$ and for $1\leq i\leq n$ we define $f_{i}^{\prime}\in F^{\prime}$ by
$f_{i}^{\prime}\left(\sum_{j=1}^{n}c_{j}e_{j}\right)=c_{i},$ (6)
and extend $f_{i}^{\prime}$ to all of $E_{r}$ by the Hahn-Banach Theorem. Then we see that $\tilde{\chi}_{ij}(s)=f_{j}^{\prime}(\pi(s)e_{i})$, which yields the desired result.
We will now show that $\pi_{r}$ is weakly contained in $U$. Let $(e_{j})_{j=1}^{n}$ be a finite unitary system in $E_{r}$ and let $f^{\prime}\in E^{\prime}$. We see that
$f^{\prime}\left(\pi(s)e_{i}\right)=\sum_{j=1}^{n}\chi_{ij}(\pi(s))f^{\prime}(e_{j}),$ (7)
so the matrix coefficient $s\mapsto f^{\prime}(\pi(s)e_{i})$ is a linear combination of matrix coefficients of $U$. Equality (5) now yields the desired result.
Now let us consider the case in which $E$ is a Hilbert space. Let $\mathscr{S}$ be the closure of $\pi(S)$ in the weak topology of $\mathcal{L}(E)$, let $\mathscr{S}_{r}$ denote the restriction of $\mathscr{S}$ to $E_{r}$, and observe that $\mathscr{S}_{r}$ is the weak closure of $\pi_{r}(S)$. We have already seen that $\mathscr{S}_{r}$ is a compact (hence amenable) topological group. Day [15] showed that any uniformly bounded representation of an amenable group $G$ on a Hilbert space is unitarizable, i.e., is equivalent to a unitary representation of $G$. Applying this to $G=\mathscr{S}_{r}$, we get that the identity map is equivalent to a unitary representation $U^{\prime}$ of $\mathscr{S}_{r}$ on a Hilbert space $\mathcal{H}$. The Peter-Weyl Theorem tells us that any unitary representation of a compact group decomposes into a direct sum of finite dimensional irreducible representations. The desired result follows from the observation that $\pi_{r}$ is equivalent to the unitary representation $U=U^{\prime}\circ\pi_{r}$. ∎